Evaluate ∫0^{π/2a}asinαxdx,α>0 by using Leibniz formula.`


1
Expert's answer
2021-08-23T04:49:42-0400
"\int_0^{π/2a}a\sin axdx=\cos ax|_0^{\pi/2a}\\=\cos \pi/2-\cos 0=0-1=-1"